### v1

# A Methodology for Improving Accuracy of Embedded Spiking Neural Networks through Kernel Size Scaling

Abstract
Spiking Neural Networks (SNNs) can offer ultra low power/ energy consumption for machine learning-based applications due to their sparse spike-based operations. Currently, most of the SNN architectures need a significantly larger model size to achieve higher accuracy, which is not suitable for resource-constrained embedded applications. Therefore, developing SNNs that can achieve high accuracy with acceptable memory footprint is highly needed. Toward this, we propose a novel methodology that improves the accuracy of SNNs through kernel size scaling. Its key steps include investigating the impact of different kernel sizes on the accuracy, devising new sets of kernel sizes, generating SNN architectures based on the selected kernel sizes, and analyzing the accuracy-memory trade-offs for SNN model selection. The experimental results show that our methodology achieves higher accuracy than state-of-the-art (93.24% accuracy for CIFAR10 and 70.84% accuracy for CIFAR100) with less than 10M parameters and up to 3.45x speed-up of searching time, thereby making it suitable for embedded applications.
Index Terms:
Spiking neural networks, kernel size scaling, embedded applications, neural architecture search.
I Introduction
Recent studies have shown that SNN algorithms can solve many machine learning (ML) applications, such as computer vision [1], robotics [2], automotive [3], and healthcare [4], while incurring ultra low power/energy consumption. Currently, the existing SNN models still consume huge model footprints to achieve high accuracy. For instance, to achieve $\geq$ 93% accuracy for CIFAR10, most of SNN models require more than 20M parameters as shown in Fig. 1(a), which is not suitable for resource-constrained embedded applications. Hence, the research problem is how can we develop SNN architectures that achieve high accuracy with acceptable memory footprint for embedded applications.
A recent work [5] has shown that developing SNN architectures in spiking domain can achieve higher accuracy than the ones derived from ANN architectures (non-spiking domain). However, this work considers limited convolution kernel sizes (i.e., 1x1 and 3x3), thus the effectiveness of its feature extraction is sub-optimal. To demonstrate this limitation, we perform experimental case study to evaluate the accuracy of SNN architectures with different sets of kernel sizes. To do this, we generate SNN architectures using neural architecture search (NAS) concept from [5] while considering two scenarios: (1) one architecture is based on 1x1 and 3x3 kernel sizes, and (2) another one is based on 1x1 and 5x5 kernel sizes. Note, the NAS concept is explained further in Section II. The experimental results are presented in Fig. 1(b), and these results indicate that employing a larger kernel size (5x5) can improve the accuracy due to its better spatial feature extraction than the smaller one (3x3). This observation exposes the opportunity to improve SNN accuracy by scaling the kernel size, while considering acceptable memory footprint for embedded applications.
Our Novel Contributions: To address the research problem and related challenges, we propose a novel methodology that improves the SNN accuracy through kernel size scaling, while considering the memory footprint. The overview of our novel methodology is shown in Fig. 2(a), and its detailed explanation is provided in Section III.
Figure 1: (a) Accuracy and memory footprints of different SNNs considering the CIFAR10 dataset: AutoSNN [6], ResNet11 [7], ResNet19 [8], CIFARNet1 [9], CIFARNet2 [10]; adapted from studies in [6]. (b) Experimental results considering the CIFAR100 dataset and two SNN architectures with different sets of kernel sizes: one architecture with 1x1 and 3x3, and another one with 1x1 and 5x5.
II Background: NAS for SNNs
NAS for SNNs is still at the early stage, hence only few studies can be found in the literature. One of the state-of-the-art is SNASNet [5], whose concept is employing a macro-architecture that has 2 neural cells; see Fig. 2(b). Each cell consists of 4 nodes and edges. Each edge refers to a specific operation from the pre-defined operations: Zeroize, SkipCon, 1x1Conv, 3x3Conv, and 3x3AvgPool for SNASNet [5], as shown by ‘black text’ in Fig. 2(c). In the searching process, different combinations of operations inside the cell are explored to find the one that can achieve high accuracy. In this work, we leverage the NAS approach from [5] with feed-forward topology to evaluate the impact of kernel size scaling and generate the SNN architectures.
III Our Proposed Methodology
Figure 2: (a) Our proposed methodology with novel contributions highlighted in blue boxes. (b) SNN macro-architecture. (c) A cell is a directed acyclic graph with each edge refers to a specific operation from the pre-defined operations. Note, our additional kernel sizes and operations are written in ‘blue text’.
Figure 3: Experimental results with CIFAR10 and CIFAR100: (a) Accuracy of different SNNs. (b) Normalized searching time of different SNNs. (3) Accuracy-memory trade-offs of different SNNs. Note, the state-of-the-art setting is 1x1_3x3_5000x, while the other settings are employed by our methodology.
Our proposed methodology employs several key steps as shown in Fig. 2(a), which are discussed in the following.
Step- 1 Investigating the impact of employing different kernel sizes on the accuracy: It aims at understanding whether employing a larger kernel size can improve accuracy. To do this, we employ NAS to generate a network architecture based on the given sets of kernel sizes. Here, we consider 1x1, 3x3, 5x5, and 7x7 kernel sizes. The 1x1 kernel aims at minimizing the computational requirements via point-wise convolution (1x1Conv) [11], while the 3x3, 5x5, and 7x7 kernels aim at extracting features through convolution operations (i.e., 3x3Conv, 5x5Conv, and 7x7Conv, respectively). Afterward, the generated architecture is trained and tested subsequently to evaluate the accuracy. The investigation results indicate that, in general, a larger kernel size can improve the accuracy as it can capture more information for unique feature extraction.
Step- 2 Devising the new sets of kernel sizes: It aims at defining the kernel sizes (and hence the respective operations) for developing a network architecture. In the state-of-the-art work [5], only 1x1Conv and 3x3Conv operations are considered, which we refer to as Set-1. In this work, we devise several new sets of operations based on the kernel sizes explored in Step- 1, i.e., Set-2: (1x1Conv, 5x5Conv), Set-3: (3x3Conv, 5x5Conv), Set-4: (1x1Conv, 7x7Conv), Set-5: (3x3Conv, 7x7Conv), and Set-6: (5x5Conv, 7x7Conv). These different sets are employed for exploring SNN architecture candidates that can offer high accuracy with acceptable memory footprint.
Step- 3 Generating SNN architectures based on the sets of kernel sizes and reduced search iteration: It aims at developing SNN architectures based on the new sets of kernel sizes (operations) with a reduced number of searching iteration. To do this, we leverage NAS concept from the state-of-the-art [5] which employs a random search mechanism based on the defined number of search iteration to determine the cell architecture in the network; see Fig. 2(b)-(c). This NAS concept also avoids the training process for evaluating the network, rather it quickly measures the representation capability score ($R$) of an investigated architecture for achieving high accuracy at the early stage through the following key steps: (1) input samples in a mini-batch are fed to the network, and the neuron activities in each layer are recorded as binary vector $f$; (2) a matrix K for each layer is formed by calculating the Hamming distance $H(f_{i},f_{j})$ between different samples $i$ and $j$; and (3) the $R$ score of the network is calculated using Eq. 1. Afterward, the network with the highest $R$ is selected as a network candidate.
$\begin{split}R=\log(\det\;\lvert\sum_{l}\textbf{K}^{l}\rvert)\;\;\;\text{with}\;l\;\text{denotes the network layer-}l\end{split}\vspace{-0.3cm}$ (1)
Each network candidate is then evaluated for its accuracy and memory size. The memory size is represented by the number of parameters ($P$), which can be calculated using Eq. 2, following the studies in [12]. $P_{w}^{i}$ and $P_{b}^{i}$ represent the number of weights and bias in layer-$l$, respectively. The dimension of weight filters in the given layer is denoted with $h_{1}$ for height, $h_{2}$ for width, $c$ for the number of channel, and $r$ for the number of filters.
$\begin{split}P=\sum_{l}(P_{w}^{l}+P_{b}^{l})\;\;\;\text{where}\;\;\;P_{w}=h_{1}\cdot h_{2}\cdot c\cdot r\end{split}\vspace{-0.5cm}$ (2)
Furthermore, our methodology employs a fewer search iteration (e.g., 1000x) than the state-of-the-art work (i.e., 5000x search iteration), to reduce the searching time while maintaining high accuracy.
Step- 4 Analyzing the accuracy-memory trade-offs for model selection: Different SNN architectures that are obtained from Step- 3 have different memory requirements, hence they might not satisfy the memory constraints of the applications. Toward this, we analyze the accuracy-memory trade-offs to select the appropriate SNN models.
IV Experimental Results and Discussion
[…]
Accuracy Improvements: Experimental results of the accuracy are shown in Fig. 3(a). For CIFAR10, accuracy scores of the networks with larger kernel sizes are comparable to the state-of-the-art, as they are within 1% accuracy range from the state-of-the-art; see 1. For instance, the network with 3x3_5x5_1000x achieves 93.24% accuracy after 300 training epoch while the state-of-the-art achieves 92.85% accuracy; see 2. For CIFAR100, accuracy scores of the networks with larger kernel sizes are significantly higher than the state-of-the-art across all training epochs; see 3. For instance, the network with 1x1_5x5_1000x achieves 70.84% accuracy after 300 training epoch while the state-of-the-art achieves 66.78% accuracy; see 4. Furthermore, we can also achieve iso-accuracy like the state-of-the-art using fewer training epochs (e.g., 60-100 epochs); see 5. The reason for all these accuracy improvements is that the larger kernel sizes can extract more unique input features (information) for generating a unique output feature. It is especially beneficial for tasks with large number of classes, as they have large number of unique features to distinguish. Hence, the CIFAR100 task can benefit more from larger kernel sizes than the CIFAR10 task as indicated by notable accuracy improvements from the-state-of-the-art across all training epochs.
Reduced Searching Time: Experimental results of the searching time are shown in Fig. 3(b). These results show that our methodology leads to a faster searching time than the state-of-the-art, since it employs a smaller number of search iteration (1000x) than the state-of-the-art (5000x), while maintaining high accuracy. For CIFAR10, the network with 3x3_5x5_1000x setting achieves about 3x searching time speed-up than the state-of-the-art; see 6. For CIFAR100, the network with 3x3_5x5_1000x setting achieves about 3.45x searching time speed-up than the state-of-the-art; see 7.
Accuracy-Memory Trade-Offs: Correlation graphs between accuracy and memory footprint are shown in Fig. 3(c). These graphs show that networks with larger kernel sizes usually offer higher accuracy than smaller ones, while incurring larger memory footprints. To find the appropriate network for each task, we apply the memory constraint on the correlation graph, and then select the one that meets this constraint, while offering high accuracy. For instance, if the application imposes 106 (or 10M) parameters as the memory budget, then we can select the network with 3x3_5x5_1000x for CIFAR10 (see 8), and the one with 1x1_5x5_1000x for CIFAR100 (see 9).
In summary, all these results show that our novel methodology can offer high accuracy with acceptable memory footprint and fast searching time, thereby enabling efficient deployment of SNNs for embedded applications.
V Conclusion
We propose a novel methodology that improves the SNN accuracy through kernel size scaling, i.e., by investigating the impact of kernel sizes on accuracy, devising the new sets of kernel sizes, generating SNN architectures based on the kernel sizes, and analyzing accuracy-memory trade-offs for model selection. Experimental results show that our methodology offers higher accuracy than state-of-the-art with acceptable memory footprint and reduced searching time, hence enabling efficient deployment of SNNs for embedded applications.
Acknowledgment
This work was partially supported by the NYUAD Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010. This work was also partially supported by the project “eDLAuto: An Automated Framework for Energy-Efficient Embedded Deep Learning in Autonomous Systems”, funded by the NYUAD Research Enhancement Fund (REF).

### v2

@@
 Acknowledgment
-This work was partially supported by the NYUAD Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.
-This work was also partially supported by the project “eDLAuto: An Automated Framework for Energy-Efficient Embedded Deep Learning in Autonomous Systems”, funded by the NYUAD Research Enhancement Fund (REF).
+This work has been supported in parts by the NYUAD Center for Cyber Security (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104, and Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.